# Three-dimensional fracton topological orders with boundary Toeplitz braiding

### v4

@@ Abstract
In this paper, we theoretically study a class of 3D non-liquid states that show exotic boundary phenomena in the thermodynamical limit. More concretely, we focus on a class of 3D fracton topological orders formed via stacking 2D twisted $\mathbb{Z}_{N}$ topologically ordered layers along $z$-direction. Nearby layers are coupled while maintaining translation symmetry along $z$ direction. The effective field theory is given by the infinite-component Chern-Simons (iCS) field theory, with an integer-valued symmetric block-tridiagonal Toeplitz $K$-matrix whose size is thermodynamically large. With open boundary conditions (OBC) along $z$, certain choice of $K$-matrices exhibits exotic boundary “Toeplitz braiding”, where the mutual braiding phase angle between two anyons at opposite boundaries oscillates and remains non-zero in the thermodynamic limit. In contrast, in trivial case, the mutual braiding phase angle decays exponentially to zero in the thermodynamical limit. As a necessary condition, this phenomenon requires the existence of boundary zero modes in the $K$-matrix spectrum under OBC. We categorize nontrivial $K$-matrices into two distinct types. Each type-I possesses two boundary zero modes, whereas each type-II possesses only one boundary zero mode. Interestingly, the integer-valued Hamiltonian matrix of the familiar 1D “Su-Schrieffer-Heeger model” can be used as a non-trivial $K$-matrix. Importantly, since large-gauge-invariance ensures integer quantized $K$-matrix entries, global symmetries are not needed to protect these zero modes. We also present numerical simulation as well as finite size scaling, further confirming the above analytical results. Symmetry fractionalization in iCS field theory is also briefly discussed. Motivated by the present field-theoretical work, it will be interesting to construct 3D lattice models for demonstrating Toeplitz braiding, which is left to future investigation.
@@ I Introduction
As the analysis of lattice models indicates that many FTOs can be constructed by properly stacking and coupling lower-dimensional topologically-ordered liquid states. This prompts the question of what the field-theoretic counterpart is for the process of stacking topological orders. Along this line of thinking, one solution is to take full use of so-called infinite-component Chern-Simons theory (iCS), which is a very useful and powerful (3+1)D field theory recently studied in Refs. [9, 10, 11, 12, 13, 14]. Let us elaborate more.
@@
In this paper, we focus on iCS field theories in which each layer is a twisted $\mathbb{Z}_{m}$ topological order described by a two-component Chern-Simons theory $\frac{1}{4\pi}\sum^{2}_{I,J=1}\mathcal{K}_{IJ}a_{\mu}^{I}\partial_{\nu}a_{\lambda}^{J}\epsilon^{\mu\nu\lambda}$. Here, $\mathcal{K}=\left(\begin{smallmatrix}n&m\\ m&0\end{smallmatrix}\right)$ with $m,n\in\mathbb{Z}$. $a_{\mu}^{I}$ is a compact $U(1)$ gauge field restricted on the $I$-th layer, meaning that $I$ may serve as the coordinate along $z$ direction (assuming that each layer is on $xy$ plane). After stacking these (2+1)D twisted topological orders and coupling the nearby layers by additional Chern-Simons terms (see Fig. 1a), we can find that the $K$-matrix of the resulting iCS field theory is a block-tridiagonal Toeplitz matrix (to be shown in Eq. (2) in Sec. II). [^2] In the following text, the term “$K$-matrix” primarily refers to the Toeplitz $K$-matrix, as represented by Eq. (2), unless explicitly stated otherwise. Locality of the resulting 3D theory requires that $K_{IJ}$ must decay no slower than exponential decay as $|I-J|$ increases. For the sake of simplicity, we only consider the coupling between nearest (2+1)D topological orders.
@@
Mathematically, Toeplitz matrices are defined as matrices possessing translation symmetry along each diagonal. They find wide applications in both physics and mathematics [30, 31, 32, 33, 34, 35, 36]. When individual entries of a Toeplitz matrix are replaced by matrix blocks, the resulting matrix is referred to as a block Toeplitz matrix. Interestingly, the formalism of the block-tridiagonal Toeplitz matrix immediately reminds us of a single-particle Hamiltonian of a 1D tight-binding lattice fermion model. Research on topological insulators (TIs) and topological superconductors (TSCs) has shed light on some Toeplitz matrices possessing eigenvectors localized at boundaries [37, 38, 39, 40, 41, 42] when open boundary condition (OBC) is imposed. These eigenvectors have eigenvalues that decrease exponentially to zero as the size of the matrix increases, which is well-known as boundary zero modes in the thermodynamical limit. We must stress that, such an analogy is made only on the level of similar mathematics. It is important to note that $K$-matrix here is not a Hamiltonian, and all matrix elements here are strictly quantized to integers due to large gauge invariance. As is known from the study of TIs and TSCs, boundary zero modes contain rich topological physics. Inspired by research in TIs and TSCs, we wonder whether and how the potential presence of boundary zero modes of the matrix $K$ reshape the low-lying energy physics of (3+1)D fracton topological order described by iCS field theory. Especially, in the presence of boundary zero modes of $K$-matrices, do there exist nontrivial properties on $z$-boundaries (i.e., two boundaries along $z$ direction) of iCS field theory?
@@
Figure 1: (a) is an illustrative diagram of stacking $\mathbb{Z}_{m}$ twisted topological orders. After stacking these $\mathbb{Z}_{m}$ twisted topological orders, we obtain a 3-dimensional theory as illustrated in (b). In the following discussion, open boundary condition is applied in $z$-direction. In our discussion, we focus on the asymptotic behaviour as the number of layer $N$ goes to infinity. (b) is also an illustrative diagram of braiding process between excitations residing at distinct $\bm{z}$-boundaries. The topological excitations are plotted as red dots in (b). If the $K$-matrix of the iCS field theory has boundary zero mode, together with the condition that the upper-right and lower-left entries of $K^{-1}$ are non-integers, two topological excitations residing at distinct $z$-boundaries can be detected mutually via mutual braiding (i.e., Toeplitz braiding) in the thermodynamical limit.
@@
In this paper, we investigate these questions through certain iCS field theories under OBC in the $z$-direction (denoted as $z$-OBC). Specifically, we couple twisted $\mathbb{Z}_{m}$ topological orders layer-by-layer, which leads to $K$-matrices that take the form of block Toeplitz matrices. It should be noted that, again, the “boundary zero modes” of Toeplitz $K$-matrices discussed here are not physical edge modes of TIs. Instead, they are, in a pure mathematical sense, eigenvalues of dimensionless $K$-matrices with OBC, which decrease exponentially to zero as the size of the $K$-matrix, i.e., the number of layers $N$, increases from finite to infinite. Due to the presence of boundary zero modes in the $K$-matrix, we observe that the braiding statistics between topological excitations, even when widely separated along the $z$ direction, may be nontrivial, except for some very special cases. As an extreme example, Fig. 1(b) illustrates the braiding process between topological excitations residing at distinct $z$-boundaries, which also clarifies how the braiding of two excitations is practically performed. In this paper, this mutual braiding statistics is termed “Toeplitz braiding” in order to highlight the nontrivial mathematical properties of Toeplitz matrices.
@@
More analytically, we find that the presence of Toeplitz braiding necessitates not only the existence of boundary zero modes of $K$-matrices but also non-integer values of the upper-right and lower-left components of $K^{-1}$ (to be shown in Fig. 4(b) in the main text). We classify $K$-matrices that potentially support Toeplitz braiding into two distinct types and numerically verify how the braiding phase angle constantly oscillates as the number of layers along the $z$ direction tends to infinity. To make a clear comparison, we also numerically compute a trivial case, where the mutual braiding phase angle exponentially decays to zero in the thermodynamical limit. By taking full advantage of the properties of Toeplitz matrices and making a tight connection to the calculation of braiding statistics of Chern-Simons gauge theory, our work presented here illuminates the exotic boundary topological physics of 3D strongly-correlated states, advancing the exploration of topologically ordered non-liquids including fracton topological orders. Several interesting future directions are straightforward.
@@
The paper is organized as follows. In Sec. II, we carry on an analytic analysis on a concrete iCS field theory with Toeplitz braiding, whose $K$-matrix is, in its mathematical form, equivalent to the Hamiltonian of Su-Schrieffer-Heeger model [37]. In Sec. III, we propose a general framework for sorting $K$-matrices with boundary zero modes. Within this framework, we identify two distinct types of $K$-matrices that possess boundary zero modes, providing them as illustrative examples. The results are shown in Table 1. To demonstrate the correspondence between boundary zero modes of $K$-matrices and nonlocal braiding statistics in $z$-direction in these models, we conduct numerical computations in Sec. IV. Numerical computations on symmetry fractionalization of iCS field theories with Toeplitz braiding are also briefly discussed in Sec. IV. This work is concluded in Sec. V.
@@
Table 1: A class of Toeplitz $K$-matrices (Eq. (2)) with boundary zero modes, where the block $A=\begin{pmatrix}n&m\\ m&0\end{pmatrix}$ and the block $B=\begin{pmatrix}l_{1}&l_{2}\\ l_{3}&l_{4}\end{pmatrix}$. Type-I $K$-matrix possesses 2 boundary zero mode, whereas Type-II $K$-matrix possesses 1 boundary zero modes. Considering the physical meaning of block A, we require $m\neq 0$. It should be noted that the “requirement” column is not the sufficient condition for Toeplitz braiding.
@@ II Construction of iCS field theory and Toeplitz Braiding
The construction of an iCS field theory can be regarded as the process of stacking $\mathbb{Z}_{m}$ topological orders. We initiate our analysis from the $\mathbb{Z}_{m}$ twisted topological order, which is described by the $K$-matrix Chern-Simons theory $\mathcal{L}=\frac{1}{4\pi}\sum_{I,J}A_{IJ}a_{\mu}^{I}\partial_{\nu}a_{\lambda}^{J}\epsilon^{\mu\nu\lambda}$. Here, $A$ is specified as $A=\left(\begin{smallmatrix}n&m\\ m&0\end{smallmatrix}\right)$. Here, $a_{\mu}^{I}$ is a $U(1)$ gauge field and $n,m$ are quantized to integers due to the invariance under large gauge transformations. Given a stack of $\mathbb{Z}_{m}$ twisted topological order, we add Chern-Simons terms to couple the nearby layers, which corresponds to the off-diagonal blocks in the $K$-matrix. The stacking direction is dubbed the $z$-direction hereafter, and the resulting theory is described by the following Lagrangian:
@@
with $A=\begin{pmatrix}n&m\\ m&0\end{pmatrix}$ and $B=\begin{pmatrix}l_{1}&l_{2}\\ l_{3}&l_{4}\end{pmatrix}$, as shown in Eq. (2). Fig. 1(a) gives an illustration of the stacking process. Each layer is a $\mathbb{Z}_{m}$ twisted topological order described by $K$-matrix Chern-Simons theory corresponding to diagonal blocks $A$ in the Toeplitz $K$-matrix. The nearby layers are coupled by Chern-Simons terms, which corresponds to the upper-right and lower-left $2\times 2$ blocks $B$ in the $K$-matrix. For example, Layer 1 contributes intra-layer Chern-Simons terms $\frac{1}{4\pi}[m(a_{\mu}^{1}\partial_{\nu}a_{\lambda}^{2}+a_{\mu}^{2}\partial_{\nu}a_{\lambda}^{1})+n(a_{\mu}^{1}\partial_{\nu}a_{\lambda}^{1})]\epsilon^{\mu\nu\lambda}$. The inter-layer Chern-Simons terms between Layer 1 and Layer 2 are $\frac{1}{4\pi}[l_{1}(a_{\mu}^{1}\partial_{\nu}a_{\lambda}^{3}+a_{\mu}^{3}\partial_{\nu}a_{\lambda}^{1})+l_{2}(a_{\mu}^{3}\partial_{\nu}a_{\lambda}^{2}+a_{\mu}^{2}\partial_{\nu}a_{\lambda}^{3})+l_{3}(a_{\mu}^{4}\partial_{\nu}a_{\lambda}^{1}+a_{\mu}^{1}\partial_{\nu}a_{\lambda}^{4})+l_{4}(a_{\mu}^{4}\partial_{\nu}a_{\lambda}^{2}+a_{\mu}^{2}\partial_{\nu}a_{\lambda}^{4})]\epsilon^{\mu\nu\lambda}$. If we stack $N$ layers of twisted $\mathbb{Z}_{m}$ topological order, the resulting theory is a multi-component Chern-Simons theory with a $2N$-dimensional $K$-matrix. As the number of layers $N$ goes to infinity, [^3] In this paper, taking thermodynamical limit refers to taking the limit $N\rightarrow\infty$ along $z$-direction. Each $xy$ layer is already in the thermodynamical limit. the mCS theory becomes an iCS field theory as illustrated in Fig. 1(b). All the coefficients in the $K$-matrix are integers due to large gauge invariance. In this setting, the superscript $I$ of a gauge field $a_{\mu}^{I}$ carries the meaning of the $z$-coordinate. The $I$-th layer contains two independent gauge fields $a_{\mu}^{2I-1}$ and $a_{\mu}^{2I}$. We should be extremely careful about the general basis transformation $K\rightarrow W^{T}KW,\ W\in GL(2N,\mathbb{Z})$, requiring it to preserve the locality in the $z$-direction. The absence of upper-right and lower-left blocks indicates that we are discussing iCS field theories in $z$-OBC.

### v5

@@ Abstract
In this paper, we theoretically study a class of 3D non-liquid states that show exotic boundary phenomena in the thermodynamical limit. More concretely, we focus on a class of 3D fracton topological orders formed via stacking 2D twisted $\mathbb{Z}_{N}$ topologically ordered layers along $z$-direction. Nearby layers are coupled while maintaining translation symmetry along $z$ direction. The effective field theory is given by the infinite-component Chern-Simons (iCS) field theory, with an integer-valued symmetric block-tridiagonal Toeplitz $K$-matrix whose size is thermodynamically large. With open boundary conditions (OBC) along $z$, certain choice of $K$-matrices exhibits exotic boundary “Toeplitz braiding”, where the mutual braiding phase angle between two anyons at opposite boundaries oscillates and remains non-zero in the thermodynamic limit. In contrast, in trivial case, the mutual braiding phase angle decays exponentially to zero in the thermodynamical limit. As a necessary condition, this phenomenon requires the existence of boundary zero modes in the $K$-matrix spectrum under OBC. We categorize nontrivial $K$-matrices into two distinct types. Each type-I possesses two boundary zero modes, whereas each type-II possesses only one boundary zero mode. Interestingly, the integer-valued Hamiltonian matrix of the familiar 1D “Su-Schrieffer-Heeger model” can be used as a non-trivial $K$ matrix. Importantly, since large-gauge-invariance ensures integer quantized $K$-matrix entries, global symmetries are not needed to protect these zero modes. We also present numerical simulation as well as finite size scaling, further confirming the above analytical results. Symmetry fractionalization in iCS field theory is also briefly discussed. Motivated by the present field-theoretical work, it will be interesting to construct 3D lattice models for demonstrating Toeplitz braiding, which is left to future investigation.
@@ I Introduction
As the analysis of lattice models indicates that many FTOs can be constructed by properly stacking and coupling lower-dimensional topologically-ordered liquid states. This prompts the question of what the field-theoretic counterpart is for the process of stacking topological orders. Along this line of thinking, one solution is to take full use of so-called infinite-component Chern-Simons theory (iCS), which is a very useful and powerful (3+1)D field theory recently studied in Refs. [9, 10, 11, 12, 13, 14, 15]. Let us elaborate more.
@@
In this paper, we focus on iCS field theories in which each layer is a twisted $\mathbb{Z}_{m}$ topological order described by a two-component Chern-Simons theory $\frac{1}{4\pi}\sum^{2}_{I,J=1}\mathcal{K}_{IJ}a_{\mu}^{I}\partial_{\nu}a_{\lambda}^{J}\epsilon^{\mu\nu\lambda}$. Here, $\mathcal{K}=\left(\begin{smallmatrix}n&m\\ m&0\end{smallmatrix}\right)$ with $m,n\in\mathbb{Z}$. $a_{\mu}^{I}$ is a compact $U(1)$ gauge field restricted on the $I$-th layer, meaning that $I$ may serve as the coordinate along $z$ direction (assuming that each layer is on $xy$ plane). After stacking these (2+1)D twisted topological orders and coupling the nearby layers by additional Chern-Simons terms (see Fig. 1a), we can find that the $K$ matrix of the resulting iCS field theory is a block-tridiagonal Toeplitz matrix (to be shown in Eq. (2) in Sec. II). [^2] In the following text, the term “$K$ matrix” primarily refers to the Toeplitz $K$ matrix, as represented by Eq. (2), unless explicitly stated otherwise. Locality of the resulting 3D theory requires that $K_{IJ}$ must decay no slower than exponential decay as $|I-J|$ increases. For the sake of simplicity, we only consider the coupling between nearest (2+1)D topological orders.
@@
Mathematically, Toeplitz matrices are defined as matrices possessing translation symmetry along each diagonal. They find wide applications in both physics and mathematics [31, 32, 33, 34, 35, 36, 37]. When individual entries of a Toeplitz matrix are replaced by matrix blocks, the resulting matrix is referred to as a block Toeplitz matrix. Interestingly, the formalism of the block-tridiagonal Toeplitz matrix immediately reminds us of a single-particle Hamiltonian of a 1D tight-binding lattice fermion model. Research on topological insulators (TIs) and topological superconductors (TSCs) has shed light on some Toeplitz matrices possessing eigenvectors localized at boundaries [38, 39, 40, 41, 42, 43] when open boundary condition (OBC) is imposed. These eigenvectors have eigenvalues that decrease exponentially to zero as the size of the matrix increases, which is well-known as boundary zero modes in the thermodynamical limit. We must stress that, such an analogy is made only on the level of similar mathematics. It is important to note that $K$ matrix here is not a Hamiltonian, and all matrix elements here are strictly quantized to integers due to large gauge invariance. As is known from the study of TIs and TSCs, boundary zero modes contain rich topological physics. Inspired by research in TIs and TSCs, we wonder whether and how the potential presence of boundary zero modes of the matrix $K$ reshape the low-lying energy physics of (3+1)D fracton topological order described by iCS field theory. Especially, in the presence of boundary zero modes of $K$ matrices, do there exist nontrivial properties on $z$-boundaries (i.e., two boundaries along $z$ direction) of iCS field theory?
@@
Figure 1: (a) is an illustrative diagram of stacking $\mathbb{Z}_{m}$ twisted topological orders. After stacking these $\mathbb{Z}_{m}$ twisted topological orders, we obtain a 3-dimensional theory as illustrated in (b). In the following discussion, open boundary condition is applied in $z$-direction. In our discussion, we focus on the asymptotic behaviour as the number of layer $N$ goes to infinity. (b) is also an illustrative diagram of braiding process between excitations residing at distinct $\bm{z}$-boundaries. The topological excitations are plotted as red dots in (b). If the $K$ matrix of the iCS field theory has boundary zero mode, together with the condition that the upper-right and lower-left entries of $K^{-1}$ are non-integers, two topological excitations residing at distinct $z$-boundaries can be detected mutually via mutual braiding (i.e., Toeplitz braiding) in the thermodynamical limit.
@@
In this paper, we investigate these questions through certain iCS field theories under OBC in the $z$-direction (denoted as $z$-OBC). Specifically, we couple twisted $\mathbb{Z}_{m}$ topological orders layer-by-layer, which leads to $K$ matrices that take the form of block Toeplitz matrices. It should be noted that, again, the “boundary zero modes” of Toeplitz $K$ matrices discussed here are not physical edge modes of TIs. Instead, they are, in a pure mathematical sense, eigenvalues of dimensionless $K$ matrices with OBC, which decrease exponentially to zero as the size of the $K$ matrix, i.e., the number of layers $N$, increases from finite to infinite. Due to the presence of boundary zero modes in the $K$ matrix, we observe that the braiding statistics between topological excitations, even when widely separated along the $z$ direction, may be nontrivial, except for some very special cases. As an extreme example, Fig. 1(b) illustrates the braiding process between topological excitations residing at distinct $z$-boundaries, which also clarifies how the braiding of two excitations is practically performed. In this paper, this mutual braiding statistics is termed “Toeplitz braiding” in order to highlight the nontrivial mathematical properties of Toeplitz matrices.
@@
More analytically, we find that the presence of Toeplitz braiding necessitates not only the existence of boundary zero modes of $K$ matrices but also non-integer values of the upper-right and lower-left components of $K^{-1}$ (to be shown in Fig. 4(b) in the main text). We classify $K$ matrices that potentially support Toeplitz braiding into two distinct types and numerically verify how the braiding phase angle constantly oscillates as the number of layers along the $z$ direction tends to infinity. To make a clear comparison, we also numerically compute a trivial case, where the mutual braiding phase angle exponentially decays to zero in the thermodynamical limit. By taking full advantage of the properties of Toeplitz matrices and making a tight connection to the calculation of braiding statistics of Chern-Simons gauge theory, our work presented here illuminates the exotic boundary topological physics of 3D strongly-correlated states, advancing the exploration of topologically ordered non-liquids including fracton topological orders. Several interesting future directions are straightforward.
@@
The paper is organized as follows. In Sec. II, we carry on an analytic analysis on a concrete iCS field theory with Toeplitz braiding, whose $K$ matrix is, in its mathematical form, equivalent to the Hamiltonian of Su-Schrieffer-Heeger model [38]. In Sec. III, we propose a general framework for sorting $K$ matrices with boundary zero modes. Within this framework, we identify two distinct types of $K$ matrices that possess boundary zero modes, providing them as illustrative examples. The results are shown in Table 1. To demonstrate the correspondence between boundary zero modes of $K$ matrices and nonlocal braiding statistics in $z$-direction in these models, we conduct numerical computations in Sec. IV. Numerical computations on symmetry fractionalization of iCS field theories with Toeplitz braiding are also briefly discussed in Sec. IV. This work is concluded in Sec. V.
@@
Table 1: A class of Toeplitz $K$ matrices (Eq. (2)) with boundary zero modes, where the block $A=\begin{pmatrix}n&m\\ m&0\end{pmatrix}$ and the block $B=\begin{pmatrix}l_{1}&l_{2}\\ l_{3}&l_{4}\end{pmatrix}$. Type-I $K$ matrix possesses 2 boundary zero modes, whereas Type-II $K$ matrix possesses 1 boundary zero mode. Considering the physical meaning of block A, we require $m\neq 0$. It should be noted that the “requirement” column is not the sufficient condition for Toeplitz braiding.
@@ II Construction of iCS field theory and Toeplitz Braiding
The construction of an iCS field theory can be regarded as the process of stacking $\mathbb{Z}_{m}$ topological orders. We initiate our analysis from the $\mathbb{Z}_{m}$ twisted topological order, which is described by the $K$-matrix Chern-Simons theory $\mathcal{L}=\frac{1}{4\pi}\sum_{I,J}A_{IJ}a_{\mu}^{I}\partial_{\nu}a_{\lambda}^{J}\epsilon^{\mu\nu\lambda}$. Here, $A$ is specified as $A=\left(\begin{smallmatrix}n&m\\ m&0\end{smallmatrix}\right)$. Here, $a_{\mu}^{I}$ is a $U(1)$ gauge field and $n,m$ are quantized to integers due to the invariance under large gauge transformations. Given a stack of $\mathbb{Z}_{m}$ twisted topological order, we add Chern-Simons terms to couple the nearby layers, which corresponds to the off-diagonal blocks in the $K$ matrix. The stacking direction is dubbed the $z$-direction hereafter, and the resulting theory is described by the following Lagrangian:
@@
with $A=\begin{pmatrix}n&m\\ m&0\end{pmatrix}$ and $B=\begin{pmatrix}l_{1}&l_{2}\\ l_{3}&l_{4}\end{pmatrix}$, as shown in Eq. (2). Fig. 1(a) gives an illustration of the stacking process. Each layer is a $\mathbb{Z}_{m}$ twisted topological order described by $K$-matrix Chern-Simons theory corresponding to diagonal blocks $A$ in the Toeplitz $K$ matrix. The nearby layers are coupled by Chern-Simons terms, which corresponds to the upper-right and lower-left $2\times 2$ blocks $B$ in the $K$ matrix. For example, Layer 1 contributes intra-layer Chern-Simons terms $\frac{1}{4\pi}[m(a_{\mu}^{1}\partial_{\nu}a_{\lambda}^{2}+a_{\mu}^{2}\partial_{\nu}a_{\lambda}^{1})+n(a_{\mu}^{1}\partial_{\nu}a_{\lambda}^{1})]\epsilon^{\mu\nu\lambda}$. The inter-layer Chern-Simons terms between Layer 1 and Layer 2 are $\frac{1}{4\pi}[l_{1}(a_{\mu}^{1}\partial_{\nu}a_{\lambda}^{3}+a_{\mu}^{3}\partial_{\nu}a_{\lambda}^{1})+l_{2}(a_{\mu}^{3}\partial_{\nu}a_{\lambda}^{2}+a_{\mu}^{2}\partial_{\nu}a_{\lambda}^{3})+l_{3}(a_{\mu}^{4}\partial_{\nu}a_{\lambda}^{1}+a_{\mu}^{1}\partial_{\nu}a_{\lambda}^{4})+l_{4}(a_{\mu}^{4}\partial_{\nu}a_{\lambda}^{2}+a_{\mu}^{2}\partial_{\nu}a_{\lambda}^{4})]\epsilon^{\mu\nu\lambda}$. If we stack $N$ layers of twisted $\mathbb{Z}_{m}$ topological order, the resulting theory is a multi-component Chern-Simons theory with a $2N$-dimensional $K$ matrix. As the number of layers $N$ goes to infinity, [^3] In this paper, taking thermodynamical limit refers to taking the limit $N\rightarrow\infty$ along $z$-direction. Each $xy$ layer is already in the thermodynamical limit. the mCS theory becomes an iCS field theory as illustrated in Fig. 1(b). All the coefficients in the $K$ matrix are integers due to large gauge invariance. In this setting, the superscript $I$ of a gauge field $a_{\mu}^{I}$ carries the meaning of the $z$-coordinate. The $I$-th layer contains two independent gauge fields $a_{\mu}^{2I-1}$ and $a_{\mu}^{2I}$. We should be extremely careful about the general basis transformation $K\rightarrow W^{T}KW,\ W\in GL(2N,\mathbb{Z})$, requiring it to preserve the locality in the $z$-direction. The absence of upper-right and lower-left blocks indicates that we are discussing iCS field theories in $z$-OBC.